OpenStudy (anonymous):

please will someone help me! everyones been ignoring me:( 2/6x+4/3x^2

OpenStudy (diyadiya):

What do you have to do?

OpenStudy (anonymous):

Add and simplify if possible:

OpenStudy (diyadiya):

Find the LCM Of 6x & 3x^2

OpenStudy (campbell_st):

you need a common denominator of 18x^3 \[2/6x \times 3x^2/3x^2 + 4/3x^2 \times 6x/6x = (6x^2 + 24x)/18x^3\] simplifying the fraction taking 6x as a common factor \[[6x(x + 4)]/(6x(3x^2))= (x+4)/3x^2\]

OpenStudy (diyadiya):

 So here 6x^2 is the common denominator \[\frac{2 \times x}{6x \times x^2}+ \frac{4 \times 2}{3x^2 \times 2}= \frac{2x+8}{6x^2}\] Taking common (2) out \[\frac{2(x+4)}{2(3x^2)}= \frac{x+4}{3x^2}\]